OpenStudy (anonymous):

two trains are 416 miles apart, and their speeds differ by 20 mph. They are traveling toward each other and will meet in 4 hours. Find the speed of the slower train.

OpenStudy (amistre64):

we just did this but with airplanes .... right?

OpenStudy (anonymous):

Yeah... lol but I still don't know how to exactly do it lol.. :)

OpenStudy (amistre64):

how do we measure speed?

OpenStudy (anonymous):

i think she left

OpenStudy (anonymous):

I'm here lol

OpenStudy (anonymous):

Idk how to measure the speed

OpenStudy (amistre64):

wave your arms so we can see you :)

OpenStudy (amistre64):

speed is distance divided by time

OpenStudy (anonymous):

I did... Do you see me?

OpenStudy (anonymous):

Okay.... So how do I set it up? :/

OpenStudy (amistre64):

if you travel 15 miles in 5 hours; you are going 15/5 miles per hour; or simple 5 mph

OpenStudy (amistre64):

we know the total distance; so the sum of the distances traveled by each have to equal the total right? since they meet when they meet

OpenStudy (anonymous):

true....

OpenStudy (amistre64):

Distance A = some speed times 4 hours Distance B = (some speed + 20) times 4 hours and A + B = 416

OpenStudy (anonymous):

20x=416?

OpenStudy (amistre64):

youve got the right idea, but we have to match the information in the problem :)

OpenStudy (amistre64):

416 = 4s + 4(s+20)

OpenStudy (anonymous):

amistre64 can you help with my question next plz?

OpenStudy (amistre64):

416 = 4s + 4(s+20) /4 /4 /4 ------------------- 104 = s + s + 20 104 = 2s + 20 /2 /2 -------------- 52 = s + 10 -10 -10 ------------ 42 = s

OpenStudy (anonymous):

I got 42!!!

OpenStudy (amistre64):

42 is good :)

OpenStudy (anonymous):

Awe... you beat me lol :/

OpenStudy (anonymous):

:) yay!

OpenStudy (amistre64):

by the way, not that I need any medals, some people like the notion of getting a medal for a good answer. If you dont see a "good medal" button next to a persons name you mihgt need to refresh your browser by hitting f5.